Calculate the theoretical density for Mgo. Show all your steps and draw the unit cell....
Answer #1

Unit cell of MgO is shown below

Mg + 0000 02-

Theoreticaldensity ion weight in unitcell volume of unit cell

Pth NaMa + NcMc avogadro number * V

Where

Na= number of anion per unit cell(O2-)

Ma= mass if anion = 16 g/mol

Nc = number of cation per unit cell(Mg2+)

Mc = mass 0f cation = 24.305 g/mol

Let us find Na and Nb

Each anion occupy every corner of unit cell (8 corner) and every face center (6 faces)

So

Na = 8*=+6* =1+3=4 2

Each cation occupy edge center (12 edges) and centre of unit cell.

1 Nc= 12 * =+1=3+1=4 4

Now we need to find edge length (a)

By analysing unit cell we get to know

So

a = det da

Where de, da are diameter of cation and anion

a=det da dea 172 pm da 2.52pm detda 2 252+172 2 424 pm Volume of unut cell raas (424X106033 7.6225x10 - 23 Theoretical Densit

Mg + 0000 02-

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image